OpenStudy (anonymous):

if sinx = 4/5 and x is in the first quadrant, find sin2x What I know : sin2x = 2sinxcosx (in my notes) cos2x = 1-2sin^2x = -7/25 (I solved that myself) sinx = 4/5 (given)

OpenStudy (anonymous):

Ok we now sin^2(x)+cos^2(x)=1 replace for sin(x) and find cos(x) then find sin(2x)

OpenStudy (anonymous):

My notes are telling me to use the unit circle and to use the terminal point

OpenStudy (anonymous):

Bro thats straight forward you got 16/25 +cos^2(x)=1 then cos^2(x)=9/25 then cos(x)=3/5 therfore sin(2x)=2(3/5)(4/5)=24/25 thats it i gave you the answer

OpenStudy (anonymous):

I'm a girl, not a "bro" and obviously it's not straight forward to me since I'm asking for help . I don't understand the terminal points and I'm taking an algebra 2 class and it's apparently trigonometry is apart of algebra 2

OpenStudy (anonymous):

Sorry i had no idea what sex are you since your username didn't signify anything well yeah thats trig and it seems that you wrote in your notes everything that i just used what is the problem?

OpenStudy (anonymous):

They're my initials and I don't understand how to find a terminal point . I don't understand trigonometry whatsoever

OpenStudy (anonymous):

How old are you?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

I don't see why my age is relevant lol

OpenStudy (anonymous):

Nothing its just that I've taken trig form when i was 14 😂😜 im 17 now

OpenStudy (anonymous):

If you want me I can explain trig to you but it would take time

OpenStudy (anonymous):

how do you understand this

OpenStudy (anonymous):

What trig?

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

By using scalar product in the trig circle

OpenStudy (anonymous):

what's scalar product

OpenStudy (anonymous):

Product of vectors, havent you heard of it?

OpenStudy (anonymous):

no ..

OpenStudy (anonymous):

Have you heard about the trig circle?

OpenStudy (anonymous):

is that unit circle ?

OpenStudy (anonymous):

?